# Positional Encoding in Transformer-Based Time Series Models: A Survey

### v1

@@ Introduction
Self-Attention Mechanism
Figure 1: Scaled Dot-Product Attention and Multi-Head Attention mechanisms. On the left, the Scaled Dot-Product Attention computes attention scores by scaling the dot product of query and key matrices and applying a softmax function, which is then used to weight the value matrix. On the right, the Multi-Head Attention mechanism processes multiple scaled dot-product attention operations in parallel, concatenates their outputs, and passes them through a final linear layer to generate the output. This allows the model to capture diverse contextual relationships in the data.
@@
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{x}\times d_{k}}$ are learnable weight matrices, and $d_{k}$ is the dimension of the query and key vectors.
Next, the attention scores are computed by taking the scaled dot product between the query and key matrices:
@@
where $\frac{1}{\sqrt{d_{k}}}$ is a scaling factor to prevent the dot product from growing too large, and the softmax function ensures that the attention scores sum to 1.
@@
To enhance the model’s capacity, self-attention can be computed in multiple parallel attention heads. Each head learns a distinct representation by using different sets of projection matrices:
$MultiHead(Q,K,V)=$
$Concat(head_{1},head_{2},\dots,head_{h})W_{O},$
where each $head_{i}=Attention(Q_{i},K_{i},V_{i})$, $h$ is the number of attention heads, and $W_{O}\in\mathbb{R}^{hd_{k}\times d_{x}}$ is a learnable output projection matrix.
@@
Absolute Positional Encoding
Absolute positional encoding assigns a unique positional vector to each time step in the sequence. The most common approach is using sinusoidal functions, introduced by Vaswani et al. [23]. This method allows the model to embed positional information without requiring it to learn these representations during training, making it computationally efficient for long sequences.
@@ 3.3.1 Fixed Absolute Positional Encoding
where $pos$ is the position (time step) in the sequence, $i$ is the dimension, and $d$ is the dimensionality of the positional encoding.
@@ 3.3.2 Learnable Positional Encoding
where $W_{pos}$ is an $d_{\text{model}}$-dimensional vector representing the learnable positional encoding parameters for position $pos$.
@@
$\text{Attention}(Q,K,V)=$
$\text{softmax}\left(\frac{(Q+PE)\cdot(K+PE)^{T}}{\sqrt{d_{k}}}\right)V$
@@
Relative Positional Encoding
In relative positional encoding, the attention mechanism is adjusted to incorporate the relative positions between time steps, making it more flexible in capturing dependencies between time steps regardless of their absolute positions in the sequence.
The attention score with relative positional encoding is computed as:
@@
where $x_{i}$ and $x_{j}$ are the input representations at positions $i$ and $j$, $W_{Q}$ and $W_{K}$ are the learnable projection matrices for queries and keys, and $a^{K}_{ij}$ is the relative positional encoding defined as $a^{K}_{ij}=W_{K}^{r}r_{i-j}$, where $r_{i-j}$ is the learnable relative positional embedding for the distance $i-j$.
@@
Figure 2: Comparison of Absolute and Relative Positional Encoding in Self-Attention. Absolute positional encoding is typically applied by adding fixed positional vectors to the input embeddings and modifying the queries and keys. In contrast, relative positional encoding directly incorporates relative positional information into the attention score calculation. The function $A$ represents the attention mechanism, where absolute modifies the input embeddings, while relative positional encoding adjusts the attention scores to capture positional dependencies more effectively.
As illustrated in Fig. 3, positional encoding methods can be combined to leverage the advantages of both absolute and relative approaches. This combined architecture integrates absolute positional information through additive embeddings while incorporating relative position information in the attention computation.
Time Absolute Position Encoding (tAPE)
The time Absolute Position Encoding (tAPE) improves upon the traditional sinusoidal-based absolute position encoding, which was originally designed for language models. In high-dimensional embeddings, such as 512 or 1024, the dot product between sinusoidal positional embeddings at different positions decays smoothly with increasing distance, maintaining what is referred to as distance awareness. However, for lower-dimensional embeddings (e.g., 64 or 128), this property diminishes, leading to irregularities in capturing positional relationships.
In the original transformer, the frequency term is defined as:
$\omega_{k}=10000^{-2k/d_{model}}$
tAPE modifies this by introducing sequence length consideration:
@@
where:
$d_{model}$ is the embedding dimension, $L$ is the sequence length, $\omega_{k}$ is the original frequency term and $\omega_{k}{new}$ is the adjusted frequency term
The positional encoding is then computed as:
@@
tAPE introduces a modification that incorporates the sequence length into the frequency terms of the sine and cosine functions. This ensures a smoother and more regular dot product between positional embeddings, regardless of sequence length. By adjusting the frequency scaling with the embedding dimension, tAPE retains distance awareness even for lower-dimensional embeddings, making it more effective for time series tasks.
Efficient Relative Position Encoding (eRPE)
The efficient Relative Position Encoding (eRPE) modifies traditional relative position encodings by adding the relative positional bias after the softmax operation in the attention mechanism. This approach sharpens the positional embeddings, making them more effective for time series classification (TSC), where capturing key relative positions is crucial for identifying temporal patterns.
The attention score in eRPE is computed as:
@@
$w$ is a learnable parameter vector of size $12L-$ 1 $i-j+L$ indexes the relative position between positions $i$ and $j$ $L$ is the maximum sequence length The relative position bias $w_{i-j+L}$ is added after the softmax operation
To implement eRPE efficiently, for a sequence length $L$, a trainable parameter $w$ of size $12L-$ 1 is created, representing the possible distances between positions. The relative scalar $w_{i-j+L}$ for two positions $i$ and $j$ is added to the attention scores post-softmax, ensuring sharper position distinctions. This method also optimizes GPU usage through efficient indexing operations, making it scalable for longer time series.
Transformer with Untied Positional Encoding (TUPE)
@@
Convolutional Sinusoidal Positional Encoding (ConvSPE)
@@
Temporal Positional Encoding (T-PE)
The Temporal Positional Encoding (T-PE) method combines two components to capture both positional and temporal dependencies in time series data. The geometric component applies enhanced sinusoidal positional encoding across multiple layers of the transformer to maintain positional information throughout the network:
$PE_{i,2k}=\sin\left(\frac{i}{10000^{\frac{2k}{d}}}\right),$
$\quad PE_{i,2k+1}=\cos\left(\frac{i}{10000^{\frac{2k}{d}}}\right)$
The semantic component adjusts attention scores by incorporating token similarity between positions $i$ and $j$:
@@
The final positional encoding is a combination of both components:
$T\text{-}PE(i)=PE(i)+S(i,j)$
This combined approach enables the model to capture both global and local temporal dependencies, improving performance in time series tasks.
Figure 3: A combined approach to positional encoding: Absolute + Relative. This is one method to integrate both absolute and relative positional encodings, aiming to capture both global and local positional dependencies within the attention mechanism.
Experimental Setup
The experimental setup involves comparing different positional encoding methods in two distinct Transformer-based architectures for time series data. The Time Series Transformer [29], shown in Figure 4(a), processes the raw multivariate time series data directly. For a time series with $L$ timesteps and $d$ channels, it treats each timestep as a token, resulting in $L$ tokens of dimension $d$ that are fed into the transformer layers. This architecture incorporates normalization layers after both the attention and feed-forward components to stabilize training and address internal covariate shifts. In contrast, the patch embedding transformer [2], illustrated in Figure 4(b), first transforms the input data using convolutional operations to create overlapping patches. Each patch spans multiple timesteps but maintains the full channel dimension, effectively reducing the sequence length while enriching the feature representation of each token. This results in $L/p$ tokens (where $p$ is the patch size) with increased feature dimensionality before being processed by the transformer layers. The detailed algorithms for both architectures are presented in Algorithm 1 and Algorithm 2, respectively. These architectures were chosen to evaluate positional encoding methods under different sequence processing paradigms: direct temporal processing in the Time Series Transformer versus hierarchical feature extraction through patch embedding.
@@
1
Input: Time series data $X\in\mathbb{R}^{N\times D}$
Output: Class probabilities $y\in\mathbb{R}^{C}$
2
3TimeSeriesInputEncoder:
4     Linear Projection: Map input time series $X$ to embeddings $E\in\mathbb{R}^{N\times D_{e}}$.
5     BatchNorm1d: Normalize embeddings across the batch.
6     Positional Encoding: Inject positional information into embeddings $E$.
7
8TransformerEncoder (4 Layers):
9 for $l=1$ to $4$ do
10           Layer $l$:
11              Multi-Head Attention: Compute attention-weighted representations.
12              BatchNorm1d: Normalize outputs of attention along the feature dimension.
13              Feed-Forward Network: Transform representations with a fully connected network.
14
15             BatchNorm1d: Normalize outputs of the FFN.
16
17
18Linear (Classifier):
    Output final class probabilities $y$.
Algorithm 1 Time Series Transformer
1
Input: Time series data $X\in\mathbb{R}^{N\times D}$
Output: Class probabilities $y\in\mathbb{R}^{C}$
2
3TimeSeriesPatchEmbedding:
4     Conv1d Layer: Convert input time series $X$ into patches $P\in\mathbb{R}^{N_{p}\times D_{p}}$.
5     Positional Encoding: Inject positional information into patches $P$.
6
7TransformerEncoder (4 Layers):
8 for $l=1$ to $4$ do
9           Layer $l$:
10              Multi-Head Attention: Compute attention-weighted representations.
11              Feed-Forward Network: Transform representations with a fully connected network.
12
13
14Linear (Feed-Forward):
15     Transform encoded features to a higher-dimensional space.
16
17Linear (Classifier):
    Output final class probabilities $y$.
Algorithm 2 Transformer with Patch Embedding
System Configuration
@@ 9           Layer $l$:
Dataset Train Size Test Size Length Classes Channels Type
@@
Evaluation Procedure
The experiments are designed to assess how well each positional encoding technique captures temporal dependencies and patterns in the data and how it affects the overall model performance on each architecture. The evaluation procedure includes:
@@
Results
Our experimental evaluation encompasses eight distinct positional encoding methods tested across eleven diverse time series datasets using two transformer architectures. The methods range from traditional approaches (PE, LPE) to relative encodings (RPE, eRPE) and advanced techniques (T-PE, SPE, TUPE), evaluated on both time series transformer and transformer with patch embedding architectures. The classification accuracies are presented in Tables 4 and 5, with performance visualizations provided through heatmaps in Figures 6(b) and 6(a).
Overall Performance
The experimental results demonstrate a consistent superiority of advanced positional encoding methods across both architectural configurations. SPE achieves the highest average ranking (1.727 for time series transformer, 2.090 for patch embedding), closely followed by TUPE (1.909 and 2.272) and T-PE (2.636 and 2.363). This performance hierarchy remains remarkably stable across both architectures, though the patch embedding architecture exhibits more balanced performance among the top-performing methods, as shown in Figures 5(a) and 5(b).
Sequence Length Analysis
The effectiveness of positional encoding methods shows a strong correlation with sequence length. For sequences exceeding 100 time steps, such as SelfRegulationSCP1 (896 steps) and Sleep (178 steps), advanced methods demonstrate substantial improvements, achieving 5-6% higher accuracy compared to fixed encoding approaches. Medium-length sequences (50-100 steps) show moderate improvements of 3-4%, while shorter sequences (less than 50 steps) exhibit smaller but consistent gains of 2-3%.
Dataset-Specific Performance
@@ 5.3.2 Environmental and Sensor Data
Dimensionality and Class Complexity
@@
Table 4: Accuracy comparison of Positional Encoding Methods in Time Series Transformer
Dataset Fixed Learnable Relative tAPE eRPE T-PE SPE TUPE Sleep (EEG) 0.854 0.862 0.867 0.871 0.875 0.881 0.883 0.888 ElectricDevices (Device) 0.682 0.697 0.723 0.718 0.731 0.741 0.755 0.743 FaceDetection (EEG) 0.627 0.654 0.672 0.633 0.686 0.694 0.698 0.705 MelbournePedestrian (Traffic) 0.675 0.704 0.729 0.701 0.742 0.748 0.754 0.762 SharePriceIncrease (Finance) 0.725 0.748 0.739 0.744 0.772 0.779 0.785 0.778 LSST (Other) 0.581 0.596 0.613 0.609 0.632 0.629 0.628 0.623 RacketSports (HAR) 0.742 0.762 0.775 0.761 0.804 0.811 0.813 0.819 SelfRegulationSCP1 (EEG) 0.849 0.857 0.871 0.853 0.891 0.898 0.901 0.907 UniMib (HAR) 0.841 0.855 0.865 0.852 0.876 0.882 0.879 0.878 RoomOccupancy (Sensors) 0.912 0.922 0.931 0.928 0.942 0.947 0.951 0.948 EMGGestures (EMG) 0.713 0.709 0.735 0.721 0.751 0.763 0.771 0.766 Avg Rank 7.909 6.454 5.272 6.363 3.727 2.636 1.727 1.909
Table 5: Accuracy comparison of Positional Encoding Methods in Time Series Transformer with Patch Embedding
Dataset Fixed Learnable Relative tAPE eRPE T-PE SPE TUPE Sleep (EEG) 0.841 0.852 0.858 0.851 0.864 0.872 0.876 0.879 ElectricDevices (Device) 0.673 0.694 0.761 0.693 0.762 0.763 0.811 0.779 FaceDetection (EEG) 0.618 0.642 0.671 0.653 0.678 0.686 0.674 0.695 MelbournePedestrian (Traffic) 0.668 0.702 0.724 0.682 0.733 0.742 0.753 0.745 SharePriceIncrease (Finance) 0.715 0.742 0.754 0.735 0.774 0.781 0.785 0.774 LSST (Other) 0.573 0.582 0.596 0.584 0.611 0.602 0.601 0.595 RacketSports (HAR) 0.733 0.756 0.767 0.761 0.801 0.792 0.775 0.805 SelfRegulationSCP1 (EEG) 0.836 0.844 0.829 0.842 0.856 0.871 0.883 0.875 UniMib (HAR) 0.832 0.844 0.854 0.833 0.864 0.871 0.867 0.865 RoomOccupancy (Sensors) 0.905 0.911 0.918 0.922 0.929 0.931 0.934 0.937 EMGGestures (EMG) 0.697 0.692 0.722 0.713 0.741 0.752 0.746 0.742 Avg Rank 7.818 6.454 5.272 6.363 3.363 2.363 2.090 2.272
@@
Discussion
Impact of Sequence Characteristics

### v2

@@ 3.1 Problem definition and notation
3.2 Self-Attention Mechanism
Figure 1: Scaled dot-product attention (left) and multi-head attention (right) mechanisms in transformer architecture.
@@ 3.2 Self-Attention Mechanism
followed by scaled dot-product attention:
@@
The theoretical foundations of attention mechanisms have been extensively analyzed [68, 69], revealing their universal approximation capabilities for sequence-to-sequence functions. Recent work has provided deeper insights into attention patterns and their interpretability [70], which is particularly relevant for understanding how different positional encodings influence attention weight distributions in time series contexts.
Multi-head attention concatenates parallel attention heads to learn diverse representations:
$\text{MultiHead}(Q,K,V)$ $=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h})W_{O}$
@@
The theoretical foundations of positional encoding in sequential models have been extensively analyzed from both signal processing [71, 72] and cybernetic control perspectives [73, 74].
3.3 Absolute Positional Encoding
Absolute positional encoding assigns unique vectors to each temporal position. The sinusoidal approach [10] provides computational efficiency by using predetermined trigonometric functions, eliminating the need for learned positional representations.
@@ 3.3.1 Fixed Absolute Positional Encoding
where pospos is the position (time step) in the sequence, $i$ is the dimension, and $d$ is the dimensionality of the positional encoding.
@@ 3.3.2 Learnable Positional Encoding
where $W_{pos}$ is an $d_{\text{model}}$-dimensional vector representing the learnable positional encoding parameters for position pospos.
@@
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{(Q+PE)\cdot(K+PE)^{T}}{\sqrt{d_{k}}}\right)V$
@@
3.4 Relative Positional Encoding
Shaw et al. [52] proposed modifying the attention computation to include relative position information:
@@ 3.4 Relative Positional Encoding
where $x_{i}$ and $x_{j}$ are the input representations at positions $i$ and $j$, $W_{Q}$ and $W_{K}$ are the learnable projection matrices for queries and keys, and $a^{K}_{ij}$ is the relative positional encoding defined as $a^{K}_{ij}=W_{K}^{r}r_{i-j}$, where $r_{i-j}$ is the learnable relative positional embedding for the distance iji-j.
@@
Figure 2: Absolute vs. relative positional encoding: absolute methods add position vectors to input embeddings, while relative methods directly incorporate positional relationships in attention computation.
As illustrated in Fig. 3, positional encoding methods can be combined to leverage the advantages of both absolute and relative approaches.
3.5 Time Absolute Position Encoding (tAPE)
Foumani et al. [21] identified that traditional sinusoidal encoding loses distance awareness in lower-dimensional embeddings. tAPE addresses this by incorporating sequence length into frequency computation:
@@ 3.5 Time Absolute Position Encoding (tAPE)
This modification ensures consistent distance awareness across different embedding dimensions and sequence lengths, making it particularly suitable for diverse time series applications.
3.6 Rotary Positional Encoding (RoPE)
Rotary Positional Encoding (RoPE) [66] uses rotation matrices to encode absolute positions while incorporating explicit relative position dependencies. This method applies rotations to query and key representations based on their absolute positions before computing dot-product attention:
$e_{ij}=\frac{(x_{i}W^{Q}+R_{n\theta})(x_{j}W^{K}+R_{n\theta})^{T}}{\sqrt{d_{z}}}$ (1)
where $R_{n\theta}$ is a rotation matrix that rotates by $n\theta$ radians:
$R_{n\theta}=\begin{bmatrix}\cos(n\theta)&-\sin(n\theta)\\ \sin(n\theta)&\cos(n\theta)\end{bmatrix}$ (2)
where $n$ is the position of the data, and $\theta$ is a scalar value (angle) chosen to rotate the value. RoPE combines the benefits of both absolute and relative PE, ensuring that attention calculations depend on relative distances while maintaining absolute position information. This hybrid approach has shown particular effectiveness in capturing long-range dependencies in sequential data.
3.7 Efficient Relative Position Encoding (eRPE)
This method enhances traditional relative positioning by applying bias terms after the softmax operation:
@@ 3.7 Efficient Relative Position Encoding (eRPE)
where $w\in\mathbb{R}^{2L-1}$ represents learnable relative position biases. The post-softmax application sharpens positional distinctions, making the method particularly effective for tasks requiring precise temporal localization.
3.8 Transformer with Untied Positional Encoding (TUPE)
@@
3.9 Convolutional Sinusoidal Positional Encoding (ConvSPE)
@@
3.10 Temporal Positional Encoding (T-PE)
This method combines geometric and semantic components to capture both positional and content-based temporal relationships: The semantic component adjusts attention scores by incorporating token similarity between positions $i$ and $j$:
$T\text{-}PE(i)=PE(i)+S(i,j)$
@@ 3.10 Temporal Positional Encoding (T-PE)
The geometric component maintains traditional positional information, while the semantic component $S(i,j)$ measures similarity between tokens at different positions, enabling the model to identify recurring patterns regardless of their absolute positions.
3.11 Attention with Linear Biases (ALiBi)
This method  [67] introduces a fundamentally different approach by eliminating traditional positional embeddings. Instead, ALiBi directly modifies attention scores by adding linear biases proportional to the distance between query and key positions:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+\mathbf{m}\cdot\mathbf{B}\right)V$
where $\mathbf{B}$ is the bias matrix with elements $B_{ij}=-|i-j|$ representing the absolute distance between positions $i$ and $j$, and $\mathbf{m}$ is a head-specific slope parameter controlling bias strength. The slope values are computed using a geometric sequence:
$m_{h}=2^{-\frac{8h}{H}}\text{ for }h=1,2,\ldots,H$
For time series applications, ALiBi uses symmetric distance calculations $B_{ij}=-|i-j|$ to enable bidirectional attention across temporal sequences. This parameter-free approach naturally encodes local temporal pattern preference while permitting long-range dependencies and enables extrapolation to longer sequences than seen during training.
Figure 3: The left matrix represents standard attention score computation $q_{i}\cdot k_{j}$ between queries and keys. The right matrix shows ALiBi’s linear bias values that are added to attention scores before softmax normalization.
4 Experimental Setup
The experimental setup involves comparing different positional encoding methods in two distinct Transformer-based architectures for time series data. The Time Series Transformer [20], shown in Figure 4(a), processes the raw multivariate time series data directly. For a time series with $L$ timesteps and $d$ channels, it treats each timestep as a token, resulting in $L$ tokens of dimension $d$ that are fed into the transformer layers. This architecture incorporates normalization layers after both the attention and feed-forward components to stabilize training and address internal covariate shifts. In contrast, the patch embedding transformer [49], illustrated in Figure 4(b), first transforms the input data using convolutional operations to create overlapping patches. Each patch spans multiple timesteps but maintains the full channel dimension, effectively reducing the sequence length while enriching the feature representation of each token. This results in $L/p$ tokens (where $p$ is the patch size) with increased feature dimensionality before being processed by the transformer layers. These architectures were chosen to evaluate positional encoding methods under different sequence processing paradigms: direct temporal processing in the Time Series Transformer versus hierarchical feature extraction through patch embedding.
@@ 4 Experimental Setup
4.1 System Configuration
@@ 4.2 Datasets
    The diversity of our dataset collection reflects the heterogeneous nature of time series applications across domains. Our selection criteria prioritize datasets that have been standardized in the time series classification literature [80, 81], ensuring reproducibility and enabling comparison with established baselines. Additionally, we include datasets from the Monash Time Series Forecasting Archive [82] to represent real-world forecasting scenarios, though our primary evaluation focuses on classification tasks for the reasons outlined in Section 1.
@@
Dataset Train Size Test Size Length Classes Channel Type
@@
4.3 Evaluation Procedure
Our evaluation methodology draws inspiration from comprehensive benchmarking practices established in both the NLP transformer literature [13] and time series analysis communities [16]. We employ stratified sampling and cross-validation techniques to ensure robust statistical evaluation, following guidelines established for deep learning evaluation in temporal domains [83]. The experiments are designed to assess how well each positional encoding technique captures temporal dependencies and patterns in the data and how it affects the overall model performance on each architecture. The evaluation procedure includes:
@@ 4.3 Evaluation Procedure
5 Results
Our experimental evaluation encompasses eight distinct positional encoding methods tested across eleven diverse time series datasets using two transformer architectures. The methods range from traditional approaches (PE, LPE) to relative encodings (RPE, eRPE) and advanced techniques (T-PE, SPE, TUPE), evaluated on both time series transformer and transformer with patch embedding architectures. The classification accuracies are presented in Tables 3 and 4, with performance visualizations provided through heatmaps in Figures 6(b) and 6(a).
5.1 Overall Performance
The experimental results demonstrate a consistent superiority of advanced positional encoding methods across both architectural configurations. SPE achieves the highest average ranking (1.727 for time series transformer, 2.090 for patch embedding), closely followed by TUPE (1.909 and 2.272) and T-PE (2.636 and 2.363). This performance hierarchy remains remarkably stable across both architectures, though the patch embedding architecture exhibits more balanced performance among the top-performing methods, as shown in Figures 5(a) and 5(b). These findings align with recent studies showing that advanced positional encodings provide significant benefits in complex temporal modeling tasks [84, 85].
5.2 Sequence Length Analysis
The effectiveness of positional encoding methods shows a strong correlation with sequence length. For sequences exceeding 100 time steps, such as SelfRegulationSCP1 (896 steps) and Sleep (178 steps), advanced methods demonstrate substantial improvements, achieving 5-6% higher accuracy compared to fixed encoding approaches. Medium-length sequences (50-100 steps) show moderate improvements of 3-4%, while shorter sequences (less than 50 steps) exhibit smaller but consistent gains of 2-3%. The correlation between sequence length and encoding effectiveness has been observed in other domains [86, 87], confirming the importance of sophisticated positioning strategies for extended sequences.
5.3 Dataset-Specific Performance
@@ 5.4 Architectural Impact
5.5 Dimensionality and Class Complexity
@@
Table 3: Accuracy of Positional Encoding Methods in Time Series Transformer
Dataset Absolute PE Relative PE Hybrid PE
Sin. PE Learn. PE tAPE RPE eRPE SPE ALiBi RoPE T-PE TUPE
Sleep (EEG) 0.854 0.862 0.871 0.867 0.875 0.883 0.861 0.864 0.881 0.888
ElectricDevices 0.682 0.697 0.718 0.723 0.731 0.755 0.702 0.693 0.741 0.743
FaceDetection 0.627 0.654 0.633 0.672 0.686 0.698 0.638 0.647 0.694 0.705
MelbournePedestrian 0.675 0.704 0.701 0.729 0.742 0.754 0.688 0.695 0.748 0.762
SharePriceIncrease 0.725 0.748 0.744 0.739 0.772 0.785 0.733 0.732 0.779 0.778
LSST 0.581 0.596 0.609 0.613 0.632 0.628 0.601 0.605 0.629 0.623
RacketSports 0.742 0.762 0.761 0.775 0.804 0.813 0.737 0.764 0.811 0.819
SelfRegulationSCP1 0.849 0.857 0.853 0.871 0.891 0.901 0.864 0.842 0.898 0.907
UniMib 0.841 0.855 0.852 0.865 0.876 0.879 0.853 0.859 0.882 0.878
RoomOccupancy 0.912 0.922 0.928 0.931 0.942 0.951 0.925 0.922 0.947 0.948
EMGGestures 0.713 0.709 0.721 0.735 0.751 0.771 0.705 0.715 0.763 0.766
Avg Rank 9.636 6.454 6.363 5.272 3.727 1.727 8.181 7.27 2.636 1.909
Table 4: Accuracy with Positional Encoding Methods in Transformer + Patch Embedding
Dataset Absolute PE Relative PE Hybrid PE
Sin. PE Learn. PE tAPE RPE eRPE SPE ALiBi RoPE T-PE TUPE
Sleep (EEG) 0.841 0.852 0.851 0.858 0.864 0.876 0.853 0.844 0.872 0.879
ElectricDevices 0.673 0.694 0.693 0.761 0.762 0.811 0.684 0.713 0.763 0.779
FaceDetection 0.618 0.642 0.653 0.671 0.678 0.674 0.626 0.635 0.686 0.695
MelbournePedestrian 0.668 0.702 0.682 0.724 0.733 0.753 0.672 0.690 0.742 0.745
SharePriceIncrease 0.715 0.742 0.735 0.754 0.774 0.785 0.721 0.731 0.781 0.774
LSST 0.573 0.582 0.584 0.596 0.611 0.601 0.592 0.583 0.602 0.595
RacketSports 0.733 0.756 0.761 0.767 0.801 0.775 0.722 0.749 0.792 0.805
SelfRegulationSCP1 0.836 0.844 0.842 0.829 0.856 0.883 0.849 0.831 0.871 0.875
UniMib 0.832 0.844 0.833 0.854 0.864 0.867 0.841 0.836 0.871 0.865
RoomOccupancy 0.905 0.911 0.922 0.918 0.929 0.934 0.915 0.917 0.931 0.937
EMGGestures 0.697 0.692 0.713 0.722 0.741 0.746 0.702 0.707 0.752 0.742
Avg Rank 9.636 6.454 6.363 5.272 3.363 2.090 7.818 7.363 2.363 2.272
@@
6 Discussion
6.1 Impact of Sequence Characteristics